OpenStudy (anonymous):

solve this using substitution y=0.35x+40 y=0.10x+50 Please!

OpenStudy (campbell_st):

well equate the equations 0.10x + 50 = 0.35x+ 40 subtract 0.10x from both sides 50 = 0.25x + 40 subtract 40 from both sides 10 = 0.25x I'll leave you to divide both sides of the equation by 0.25 to get the answer. Hope this helps.

OpenStudy (anonymous):

Thank you so much :))))

OpenStudy (anonymous):

putting the value of y from first equation in second equation so 0.35x+40=0.10x+50 taking x terns one side and all others on other side 0.35x-0.10x=50-40 0.25x=10 x=10/0.25 =40 :)

OpenStudy (anonymous):

thxs again you guys helped me out alot :)